# Continuous Space-Time Video Super-Resolution with 3D Fourier Fields

### v1

@@ 2 Related Work
So-called continuous STVSR (C-STVSR) represents both space and time as continuous domains and is therefore able to upsample in space and in time. It includes AVSR as well as fixed-scale STVSR (and also single-image SR) as special cases and, arguably, constitutes the most general and practical formulation. C-STVSR was pioneered by VideoINR (Chen et al., 2022), where videos are parametrized as two separate, decoupled INRs in image space and time. The temporal INR acts as an optical flow predictor, estimating the motion field from the intermediate frame at time to the keyframe as a function of , followed by backward warping to reconstruct intermediate features. However, modeling continuous backward flow fields is challenging: a location’s content – and thus its motion – varies with time. Even under linear motion, the backward flow field changes structurally over time, producing discontinuities at motion boundaries that are challenging to learn. In contrast, forward flow (key-to-intermediate) tracks each pixel’s trajectory through time as a continuous curve. For linear motion, the displacement vector field preserves its structure and simply scales with t, making it easier to model. Building on this idea, MoTIF (Chen et al., 2023) learns forward motion trajectories and applies softmax splatting (Niklaus & Liu, 2020) for forward warping, leaving occlusion handling and conflicting motions to the decoder. BF-STVSR (Kim et al., 2025) incorporates a Fourier basis in the latent space and B-splines to parametrize the motion field, but still depends on explicit warping to map the two keyframes to the intermediate. Furthermore, none of the mentioned methods has a principled mechanism for anti-aliasing. Instead, they count on implicitly learning the necessary, adaptive high-pass filtering from data, which makes the learning problem significantly more complicated and comes without any guarantees.
@@ 3.2 Videos as 3D Fourier series
The core of our representation is a collection of sinusoidal 3D bases functions in -space, parameterized by frequencies and phase shifts :
@@
exactly describes under a Gaussian PSF with variance . Aliasing-free sampling thus amounts to simply rescaling the individual basis functions by the frequency-dependent factor . Note that Eq. 4 can be implemented as a matrix multiplication followed by element-wise addition (for the shift) and scaling, which is much more efficient than explicit filtering or oversampling. Compared to existing C-STVSR methods (Chen et al., 2022; 2023; Kim et al., 2025), where a scale-appropriate PSF has to be learned from data as part of the neural model, hard-wiring the signal-theoretically correct PSF is not only more parameter-efficient, but also ensures good generalization, unaffected by training biases, cf. Becker et al. (2025). Note also, can be chosen separately for each dimension, for instance to express a Gaussian blur in space but point sampling in time, consistent with the functioning of many consumer cameras.
Contrary to other C-STVSR methods (Chen et al., 2022; 2023; Kim et al., 2025), our trigonometric expansion over can be sampled at arbitrary spatial and temporal resolutions with a single function call, without explicit frame warping. The formulation can naturally encode translational motion as simple phase shifts in the frequency domain (Kuglin, 1975), and the sinusoidal basis compactly captures frequency patterns across scales.
@@ 3.3 Conditional Fourier Parameterization
As mentioned above, in practice we use a voxel grid of local basis expansions, in line with other SR methods that do the same with local INRs (Chen et al., 2021; 2023; Kim et al., 2025). We emphasize that each grid cell still contains only a single, unified and continuous function . Global consistency across cell boundaries is maintained through the backbone encoder, which aggregates information over large spatio-temporal receptive fields to predict the local field parameters. Concretely, after obtaining a grid of semantic features from the encoder, matching the size of the input video, we apply a small convolutional network to map those features to a 3D grid of VFF parameters, , where and define the VFF at each voxel index . We find that it is not necessary to locally vary the frequencies between different grid cells. In other words, the base frequencies are learned once during training. At inference time they are kept fixed for all input videos and all cells, and only their amplitudes and phases are modulated to fit the input. Beyond saving compute, sharing a common frequency basis turns out to slightly improve stability and coherence of the reconstructed video.
@@ 3.4 Implementation and training
V3 is implemented in JAX (Bradbury et al., 2018). We use basis functions. As backbone encoder we employ RVRT (Liang et al., 2022), with the embedding dimension set to 90 and 12 attention heads. During training, the spatial upsampling factor is sampled randomly from the interval , and input data is generated with bicubic downsampling in space and subsampling of (high-speed) video frames in time, as in prior work. Training patches have pixels and 14 frames. We use standard data augmentations (random flipping, rotation, and resizing) and train for iterations with the AdamW (Loshchilov & Hutter, 2017) optimizer (lr, , , ) and a Cosine Annealing scheduler (Loshchilov & Hutter, 2016). We furthermore employ gradient clipping at a global norm of 1. As in prior work, all parameters are trained from scratch, except for the flow component in RVRT (RAFT, Teed & Deng, 2020), which we finetune only for the last iterations. We use a batch size of 16 and train on Nvidia GH200 chips. Inference only requires a single consumer GPU (Nvidia RTX 3090 Ti).
@@ 4 Experiments
We evaluate V3 on several benchmarks. Besides C-STVSR performance, we also test the special cases of spatial-only and temporal-only SR, as well as temporal consistency. Following the literature (Chen et al., 2022; 2023; Kim et al., 2025), we train on the 240 fps Adobe240 (Su et al., 2017) dataset, which consists of 133 videos at resolution 1280720 pixels. Input videos are obtained by spatial downsampling with random scaling factors and fixed temporal subsampling by a factor of to obtain input videos with 30 fps; all ground truth frames, randomly sampled in space and time serve as ground truth for supervision.
@@ 4.1 C-STVSR Performance
| Method | Vid4 |
# Par. | ||||||||
| DAINEDVR | 23.48 / 0.654 | 28.58 / 0.842 | 26.64 / 0.798 | 27.45 / 0.809 | 25.64 / 0.759 | 44.7 M | ||||||||
| ZoomingSloMo | 25.72 / 0.772 | 30.69 / 0.885 | — | 30.26 / 0.882 | — | 11.1 M | ||||||||
| TMNet | 25.96 / 0.780 | 30.14 / 0.870 | 28.83 / 0.851 | 29.41 / 0.852 | 28.30 / 0.835 | 12.3 M | ||||||||
| VideoINR | 25.61 / 0.771 | 30.26 / 0.879 | 29.41 / 0.867 | 29.92 / 0.875 | 29.27 / 0.865 | 11.3 M | ||||||||
| MoTIF | 25.79 / 0.775 | 31.04 / 0.888 | 30.04 / 0.877 | 30.63 / 0.884 | 29.82 / 0.875 | 12.6 M | ||||||||
| BF-STVSR | 25.85 / 0.777 | 31.17 / 0.890 | 30.22 / 0.880 | 30.83 / 0.888 | 30.12 / 0.880 | 13.5 M | ||||||||
| V3 | 26.76 / 0.818 | 32.93 / 0.922 | 32.24 / 0.918 | 32.85 / 0.921 | 32.24 / 0.916 | 13.7 M | ||||||||
@@
Figure 3 shows a visual comparison (on non-keyframes) for a sample from the Adobe240 dataset. V3 faithfully reconstructs high-frequency content over time. In the example, it is the only method that recovers both legible text and the characteristic “accordion” structure of the joint on the articulated bus.
@@ 4.2.1 Arbitrary-scale video SR
To measure AVSR performance, we follow the experimental settings of the dedicated AVSR literature (Li et al., 2024; Shang et al., 2024) and evaluate on the REDS (Nah et al., 2019) validation set, consisting of 240 videos with 1280720 pixels and 100 frames, captured with a GoPro camera at 24 fps. Table 2 compares V3 to other C-STVSR methods (all trained on Adobe240) in this setting. It has been reported (Shang et al., 2024) that, when operating at 1 temporal upsampling, C-STVSR methods barely outperform arbitrary-scale single-image super-resolution (AISR). We therefore include two recent examples for the latter. Indeed, V3 appears to be the first C-STVSR method to substantially surpass per-frame image AISR across all scaling factors within and outside the training distribution. We attribute the gain to our model’s better ability to transfer information between frames. The unified spatio-temporal basis gives V3 access to a larger temporal context window and lets it exploit redundancy between frames for spatial super-resolution, rather than merely avoid flickering.
@@ Appendix B Additional Quantitative Results
As a complement to Sec. 4.2.1, we also train our method specifically for the AVSR task, enabling direct comparison with prior work (Li et al., 2024; Shang et al., 2024). We denote this variant as , which is architecturally identical to V3-Large111With the number of basis functions reduced to ., but trained on the REDS training set without temporal upsampling and supervised solely on the temporal coordinates of the input frames. Table 5 reports PSNR and SSIM metrics for multiple scaling factors on the REDS validation set. Although not explicitly designed for AVSR – our model remains a more expressive C-STVSR method – clearly surpasses dedicated AVSR approaches in PSNR. For example, at scaling it exceeds the next best method by nearly 1 dB. In SSIM, it is only outperformed by ST-AVSR (Shang et al., 2024), which leverages an auxiliary pre-trained VGG16 encoder to extract multi-scale structural and textural features. It is worth noting that ST-AVSR also provides a baseline variant (B-AVSR) without this auxiliary encoder. Since likewise does not rely on external feature extractors, B-AVSR arguably offers the most direct comparison in terms of training data.
@@
Appendix C Computational Complexity
| VideoINR | MoTIF | BF-STVSR | V3 | |
| Inference time | 3.03 s | 1.88 s | 1.90 s | 1.27 s |
| VRAM | 2.6 GiB | 8.4 GiB | 10.4 GiB | 6.1 GiB |
Appendix D Use of Large Language Models

### v2

@@ 2 Related Work
So-called continuous STVSR (C-STVSR) represents both space and time as continuous domains and is therefore able to upsample in space and in time. It includes AVSR as well as fixed-scale STVSR (and also single-image SR) as special cases and, arguably, constitutes the most general and practical formulation. C-STVSR was pioneered by VideoINR (Chen et al., 2022), where videos are parametrized as two separate, decoupled INRs in image space and time. The temporal INR acts as an optical flow predictor, estimating the motion field from the intermediate frame at time to the keyframe as a function of , followed by backward warping to reconstruct intermediate features. However, modeling continuous backward flow fields is challenging: a location’s content – and thus its motion – varies with time. Even under linear motion, the backward flow field changes structurally over time, producing discontinuities at motion boundaries that are challenging to learn. In contrast, forward flow (key-to-intermediate) tracks each pixel’s trajectory through time as a continuous curve. For linear motion, the displacement vector field preserves its structure and simply scales with t, making it easier to model. Building on this idea, MoTIF (Chen et al., 2023b) learns forward motion trajectories and applies softmax splatting (Niklaus & Liu, 2020) for forward warping, leaving occlusion handling and conflicting motions to the decoder. BF-STVSR (Kim et al., 2025) incorporates a Fourier basis in the latent space and B-splines to parametrize the motion field, but still depends on explicit warping to map the two keyframes to the intermediate. Furthermore, none of the mentioned methods has a principled mechanism for anti-aliasing. Instead, they count on implicitly learning the necessary, adaptive high-pass filtering from data, which makes the learning problem significantly more complicated and comes without any guarantees.
@@ 3.2 Videos as 3D Fourier Series
The core of our representation is a collection of sinusoidal 3D basis functions in -space, parameterized by frequencies and phase shifts :
@@
exactly describes under a Gaussian PSF with variance . Aliasing-free sampling thus amounts to simply rescaling the individual basis functions by the frequency-dependent factor , where is inversely proportional to the effective sampling rate, as dictated by the Nyquist limit (see, e.g., Oppenheim et al., 1997). Note that Eq. 4 can be implemented as a matrix multiplication followed by element-wise addition (for the shift) and scaling, which is much more efficient than explicit filtering or oversampling. Compared to existing C-STVSR methods (Chen et al., 2022; 2023b; Kim et al., 2025), where a scale-appropriate PSF has to be learned from data as part of the neural model, hard-wiring the signal-theoretically correct PSF is not only more parameter-efficient, but also ensures good generalization, unaffected by training biases, cf. Becker et al. (2025). Note also, can be set to for individual dimensions, for instance to express a Gaussian blur in space but point sampling in time, or to a small constant value representing the narrow temporal PSF induced by finite exposure times in many consumer cameras. Contrary to other C-STVSR methods (Chen et al., 2022; 2023b; Kim et al., 2025), our trigonometric expansion over can be sampled at arbitrary spatial and temporal resolutions with a single function call, without explicit frame warping. The formulation can naturally encode translational motion as simple phase shifts in the frequency domain (Kuglin, 1975), and the sinusoidal basis compactly captures frequency patterns across scales.
@@ 3.3 Conditional Fourier Parameterization
As mentioned above, in practice we use a voxel grid of local basis expansions, in line with other SR methods that do the same with local INRs (Chen et al., 2021; 2023b; Kim et al., 2025). We emphasize that each grid cell still contains only a single, unified and continuous function . Global consistency across cell boundaries is maintained through the backbone encoder, which aggregates information over large spatio-temporal receptive fields to predict the local field parameters. Concretely, after obtaining a grid of semantic features from the encoder, matching the size of the input video, we apply a small convolutional network to map those features to a 3D grid of VFF parameters, , where and define the VFF at each voxel index . We find that it is not necessary to locally vary the frequencies between different grid cells. In other words, the base frequencies are learned once during training. At inference time they are kept fixed for all input videos and all cells, and only their amplitudes and phases are modulated to fit the input. Beyond saving compute, sharing a common frequency basis turns out to slightly improve stability and coherence of the reconstructed video.
@@ 3.4 Implementation and Training
V3 is implemented in JAX (Bradbury et al., 2018). We use basis functions. As backbone encoder we employ RVRT (Liang et al., 2022), with the embedding dimension set to 90 and 12 attention heads. During training, the spatial upsampling factor is sampled randomly from the interval , and input data is generated with bicubic downsampling in space and subsampling of (high-speed) video frames in time, as in prior work. Training patches have pixels and 14 frames. We use standard data augmentations (random flipping, rotation, and resizing) and train for iterations with an reconstruction loss using the AdamW (Loshchilov & Hutter, 2017) optimizer (lr, , , ) and a Cosine Annealing scheduler (Loshchilov & Hutter, 2016). We furthermore employ gradient clipping at a global norm of 1. As in prior work, all parameters are trained from scratch, except for the flow component in RVRT (RAFT, Teed & Deng, 2020), which we finetune only for the last iterations. We use a batch size of 16 and train on Nvidia GH200 chips. Inference only requires a single consumer GPU (RTX 3090 Ti).
@@ 4 Experiments
We evaluate V3 on several benchmarks. Besides C-STVSR performance, we also test the special cases of spatial-only and temporal-only SR, as well as temporal consistency. Following the literature (Chen et al., 2022; 2023b; Kim et al., 2025), we train on the 240 fps Adobe240 (Su et al., 2017) dataset, which consists of 133 videos at resolution 1280720 pixels. Input videos are obtained by spatial downsampling with random scaling factors and fixed temporal subsampling by a factor of to obtain input videos with 30 fps; all ground truth frames, randomly sampled in space and time serve as ground truth for supervision.
@@ 4.1 C-STVSR Performance
| Method | Vid4 | GoPro Center | GoPro Average | Adobe Center | Adobe Average | # Par. |
| DAINEDVR | 23.48 / 0.654 | 28.58 / 0.842 | 26.64 / 0.798 | 27.45 / 0.809 | 25.64 / 0.759 | 44.7 M |
| ZoomingSloMo | 25.72 / 0.772 | 30.69 / 0.885 | — | 30.26 / 0.882 | — | 11.1 M |
| TMNet | 25.96 / 0.780 | 30.14 / 0.870 | 28.83 / 0.851 | 29.41 / 0.852 | 28.30 / 0.835 | 12.3 M |
| VideoINR | 25.61 / 0.771 | 30.26 / 0.879 | 29.41 / 0.867 | 29.92 / 0.875 | 29.27 / 0.865 | 11.3 M |
| MoTIF | 25.79 / 0.775 | 31.04 / 0.888 | 30.04 / 0.877 | 30.63 / 0.884 | 29.82 / 0.875 | 12.6 M |
| BF-STVSR | 25.85 / 0.777 | 31.17 / 0.890 | 30.22 / 0.880 | 30.83 / 0.888 | 30.12 / 0.880 | 13.5 M |
| V3 | 26.76 / 0.818 | 32.96 / 0.923 | 32.26 / 0.919 | 32.91 / 0.922 | 32.29 / 0.917 | 13.7 M |
@@
Figure 3 shows a visual comparison (on non-keyframes) for a sample from the Adobe240 dataset. V3 faithfully reconstructs high-frequency content over time. In the example, it is the only method that recovers both legible text and the characteristic “accordion” structure of the articulated bus joint.
@@ 4.2.1 Arbitrary-scale video SR
To measure AVSR performance, we follow the experimental settings of the dedicated AVSR literature (Li et al., 2024; Shang et al., 2024) and evaluate on the REDS (Nah et al., 2019) validation set, consisting of 240 videos with 1280720 pixels and 100 frames, captured with a GoPro camera at 24 fps. Table 2 compares V3 to other C-STVSR methods (all trained on Adobe240) in this setting. It has been reported (Shang et al., 2024) that, when operating at 1 temporal upsampling, C-STVSR methods barely outperform arbitrary-scale single-image super-resolution (AISR). We therefore include three methods representative of three generations of AISR: CNN-based (LTE, Lee & Jin, 2022), modern transformer-based (CLIT, Chen et al., 2023a), and the – at the time of writing – strongest method based on hierarchical positional encoding, HIIF (Jiang et al., 2025). Indeed, V3 appears to be the first C-STVSR method to substantially surpass per-frame image AISR across all scaling factors within and outside the training distribution. We attribute the gain to our model’s better ability to transfer information between frames. The unified spatio-temporal basis gives V3 access to a larger temporal context window and lets it exploit redundancy between frames for spatial super-resolution, rather than merely avoid flickering.
@@ 4.3 Temporal Consistency
| Bicubic | VideoINR | MoTIF | BF-STVSR | V3 | V3-Large |
| 0.595 | 0.344 | 0.354 | 0.323 | 0.254 | 0.250 |
4.4 Analysis of Learned Basis Functions
Figure 6 analyzes the spatiotemporal behavior of basis functions learned by V3. The figure shows that V3 is able to pick up structure in the data, with a higher concentration of components along the coordinate axes of the polar plot (i.e., horizontal and vertical wave patterns representing axis-aligned structures). In terms of the marginal distribution of spatial frequencies, we see a non-uniform distribution with more higher frequencies than lower ones, increasing the capacity for reconstruction of fine details for super-resolution. We further visualize the average predicted magnitude (“coefficient”) of each basis function for an example sequence (Vid4, calendar), using a color map in the left subplot of Fig. 6. Magnitudes decrease with increasing frequency, consistent with results from classical Fourier analysis.
4.5 Computational Complexity
| VideoINR | MoTIF | BF-STVSR | V3 | |
| Inference time | 3.03 s | 1.88 s | 1.90 s | 1.27 s |
| VRAM | 2.6 GiB | 8.4 GiB | 10.4 GiB | 6.1 GiB |
@@ Appendix A Additional Qualitative Results
Figures 7, 8, 9, 10 and 13 show additional qualitative C-STVSR results for various datasets. Figures 11 and 12 show additional spatial- and temporal-only SR results, respectively.
@@ Appendix B Additional Quantitative Results
B.1 Comparison with AVSR Methods
As a complement to Sec. 4.2.1, we also train our method specifically for the AVSR task, enabling direct comparison with prior work (Li et al., 2024; Shang et al., 2024). We denote this variant as , which is architecturally identical to V3-Large111With the number of basis functions reduced to ., but trained on the REDS training set without temporal upsampling and supervised solely on the temporal coordinates of the input frames.
Table 6 reports PSNR and SSIM metrics for multiple scaling factors on the REDS validation set. Although not explicitly designed for AVSR – our model remains a more expressive C-STVSR method – clearly surpasses dedicated AVSR approaches in PSNR. For example, at scaling it exceeds the next best method by nearly 1 dB. In SSIM, it is only outperformed by ST-AVSR (Shang et al., 2024), which leverages an auxiliary pre-trained VGG16 encoder to extract multi-scale structural and textural features. It is worth noting that ST-AVSR also provides a baseline variant (B-AVSR) without this auxiliary encoder. Since likewise does not rely on external feature extractors, B-AVSR arguably offers the most direct comparison in terms of training data. Related qualitative samples are provided in Fig. 14.
B.2 Comparison with VSR Methods
We go one step further and compare V2.5 to standard video super-resolution (VSR) methods. In contrast to AVSR, VSR methods do not support arbitrary spatial scaling factors and are specifically trained for a single upsampling rate (usually ). In Tab. 7, we compare V2.5 with BasicVSR (Chan et al., 2021), TTVSR (Liu et al., 2022), and FTVSR (Qiu et al., 2022). We find that our method achieves the highest PSNR among all compared models and matches the best-performing approach in SSIM, despite being the only method supporting arbitrary-scale upsampling.
@@ B.2 Comparison with VSR Methods
| PSNR | SSIM | |
| BasicVSR | 27.24 | 0.825 |
| TTVSR | 27.64 | 0.840 |
| FTVSR | 27.57 | 0.839 |
| V2.5 | 27.77 | 0.840 |
B.3 Ablation Study
| Vid4 | Adobe240 | |||
| PSNR | SSIM | PSNR | SSIM | |
| 26.19 | 0.797 | 30.55 | 0.884 | |
| 26.50 | 0.809 | 31.55 | 0.903 | |
| 26.72 | 0.817 | 32.16 | 0.914 | |
| 26.74 | 0.818 | 32.29 | 0.917 | |
| Fixed | 22.85 | 0.651 | 24.80 | 0.738 |
| 50% train iters. | 26.71 | 0.817 | 32.12 | 0.914 |
| 25% train iters. | 26.60 | 0.814 | 31.84 | 0.909 |
In Tab. 8 we analyze several design choices of our method. To isolate the effect of the VFF capacity, we vary the size of the basis while keeping the rest of the model fixed. I.e., we chose different numbers of basis functions and retrain the mapping from the 256-dimensional encoder output as well as the -dimensional VFF basis.
On both Vid4 and Adobe240, a smaller basis () leads to a noticeable performance drop, which is further exacerbated when reducing to . On the other hand, increasing the dimension of the basis to only marginally improves performance. We consider a good compromise, noting that a larger basis may in some applications be warranted, and would still be supported by the current encoder capacity.
We further examine the importance of learning the VFF frequencies by pre-setting with a canonical 3D plane-wave Fourier basis. Frequency vectors are sampled uniformly from the 3D Fourier frequency space (k-space) and fixed during training. This variant performs significantly worse, which is expected: sampling 512 vectors uniformly in 3D frequency space results in a very sparse grid – only unique frequencies per axis – limiting representation power. In contrast, learning the basis functions globally allows to flexibly distribute a constrained number of basis functions so as to maximize their representational power, exploiting structure in the dataset. This is visible in Fig. 6, where learned frequencies concentrate along coordinate axes and distribute non-uniformly over space–time. Notably, learning the base incurs only a negligible overhead: just additional trainable scalars and no extra inference cost.
Finally, we ablate the number of training iterations: Reducing to 50% of iterations leads to almost no performance loss (e.g., 26.76 vs. 26.71 dB PSNR on Vid4 ). At 25% the drop is more noticeable, but V3 remains the best method by a substantial margin (e.g., V3 at 25%: 31.84 vs. second-best BF-STVSR at 100%: 30.12 dB).
B.4 Degradation Study
We analyze two types of image degradation during test time, as shown in Table 9. All experiments are conducted on the Vid4 dataset with a spatial upscaling factor of and a temporal upscaling factor of . (1) Noise only: We add zero-mean Gaussian noise with a specified standard deviation to images normalized to the range [-1, 1]. The corrupted images are then rescaled to 8-bit [0, 255] and rounded. (2) Noise + compression: We apply the same noise corruption, followed by H.264 video compression of the resulting frames. In both cases, we compare the model’s predictions against the non-degraded high-resolution ground-truth images.
| Degradation: (Gaussian noise) | |||||
| Method | |||||
| VideoINR | 25.55 / 0.766 | 25.50 / 0.763 | 25.45 / 0.759 | 25.04 / 0.729 | 24.35 / 0.677 |
| MoTIF | 25.73 / 0.771 | 25.68 / 0.729 | 25.62 / 0.763 | 25.16 / 0.730 | 24.46 / 0.675 |
| BF-STVSR | 25.82 / 0.774 | 25.77 / 0.771 | 25.71 / 0.767 | 25.28 / 0.735 | 24.53 / 0.679 |
| V3 | 26.68 / 0.814 | 26.50 / 0.806 | 26.28 / 0.796 | 25.58 / 0.757 | 25.18 / 0.721 |
| Degradation: (H.264 of noised input) | |||||
| VideoINR | 24.38 / 0.688 | 24.38 / 0.688 | 24.38 / 0.687 | 24.32 / 0.684 | 24.06 / 0.665 |
| MoTIF | 24.54 / 0.690 | 24.53 / 0.690 | 24.53 / 0.690 | 24.47 / 0.686 | 24.21 / 0.666 |
| BF-STVSR | 24.64 / 0.694 | 24.63 / 0.694 | 24.63 / 0.694 | 24.57 / 0.691 | 24.29 / 0.670 |
| V3 | 25.35 / 0.726 | 25.35 / 0.726 | 25.34 / 0.726 | 25.29 / 0.722 | 25.02 / 0.706 |
Appendix C Use of Large Language Models